OpenStudy (anonymous):

Evaluate (5 + 2)2 − 40 + 3

OpenStudy (anonymous):

(5+2)(5+2) - 40 + 3 = 25 + 10 + 10 + 4 - 40 +3. 25 + 24 - 40 + 3. 49 - 40 = 9 + 3 = 12. 12 is your final answer :)